# Deep Learning for Protein-Ligand Docking: Are We There Yet?

### v3

@@ 5 Results and discussions
In this section, we present PoseBench’s results for single and multi-ligand protein-ligand structure generation and discuss their implications for future work. Note that across all the experiments, for generative methods (or methods that use generative inputs to make their predictions), we report their performance metrics in terms of the mean and standard deviation across three independent runs of the method to gain insights into its inter-run stability and consistency. For interested readers, in Appendix C, we report the average runtime and memory usage of each baseline method to determine which methods are the most practical for real-world docking applications.
@@
Figure 2: Astex & PoseBusters dataset results for successful single-ligand docking.
@@
Figure 3: PoseBusters dataset results for successful single-ligand docking with relaxation.
5.1 Generalization to new binding pockets implies single-ligand docking performance
@@
We begin our investigations by evaluating the performance of each baseline method for single-ligand docking using the Astex Diverse and PoseBusters Benchmark datasets. Notably, for results on the PoseBusters Benchmark dataset, we perform an additional analysis where we apply post-prediction (fixed-protein) relaxation to each method’s generated ligand conformations using molecular dynamics simulations (Eastman and Pande, 2010), as originally proposed by Buttenschoen et al. (2024). Additionally, for interested readers, in Appendix G.1 we include benchmark results for flexible-protein relaxation as implemented by Lu et al. (2024). As shown in Figures 2 and 3, DiffDock-L achieves the best overall performance across the two datasets both with and without applying relaxation to its generated structures. Closely behind in performance for the PoseBusters Benchmark dataset are DynamicBind and RoseTTAFold-All-Atom following structural relaxation. Interestingly, without relaxation, AutoDock Vina combined with DiffDock-L’s predicted binding pockets achieves the second-best performance on the PoseBusters Benchmark dataset, which suggests that DiffDock-L is currently the only single-ligand deep learning method that presents a better intrinsic understanding of biomolecular physics for docking than conventional modeling tools. For interested readers, in Appendix G.2, we report e.g., pocket-only PoseBusters Benchmark experiments and RMSD violin plots for both the Astex Diverse and PoseBusters Benchmark datasets.
@@
Figure 4: CASP15 dataset results for successful multi-ligand docking with relaxation.
@@
Figure 5: CASP15 dataset results for multi-ligand PoseBusters validity rates with relaxation.
5.2 Molecule pretraining implies multi-ligand docking performance
We now turn to investigating the performance of various deep learning and conventional methods for multi-ligand docking. In Figure 4, we see that although DiffDock-L initially appears to achieve the best performance in this context, after applying structural relaxation its performance quickly diminishes. This trend holds for similar deep learning methods such as DynamicBind that were specifically trained on single-ligand protein complexes, as achieving a low RMSD for each (fragment) ligand does not rule out the existence of protein-ligand and particularly ligand-ligand steric clashes. In contrast to this trend, however, NeuralPLexer does not lose a significant fraction of accurate (fragment) ligand predictions following structural relaxation, which suggests that its multi-ligand conformations are already largely free of steric clashes. Figure 6 illustrates these steric clash trends using top-1 predictions from DiffDock-L and NeuralPLexer for CASP15 target T1187 as a case study.
Another interesting observation is that TULIP (following structural relaxation) outperforms the docking success rates of single-ligand deep learning docking methods such as DiffDock-L and DynamicBind in the context of multi-ligand docking, which suggests room for future improvement in the multi-ligand modeling capabilities of these recent deep learning baselines. To further inspect each method’s understanding of biomolecular physics for docking, in Figure 5 we report each method’s percentage of predicted complexes (whether correct or not) for which all ligand conformations in the complex are jointly considered valid according to the PoseBusters software suite (i.e., PB-Valid). In short, in the context of multi-ligands, we find that AutoDock Vina followed by DiffDock-L are tied in terms of their PoseBusters validity rates following structural relaxation, with our ensembling consensus baseline (i.e., Ensemble (Con)) as well as DynamicBind shortly behind. As an addendum, we note that NeuralPLexer’s (DynamicBind’s) predictions seem to be frequently selected by Ensemble (Con) for single (multi)-ligand targets, which suggests that NeuralPLexer (DynamicBind) produces the most consistent (i.e., similar) ligand poses for a given single (multi)-ligand protein complex. For interested readers, in Appendix G.3, we report additional results e.g., in terms of lDDT-PLI and RMSD violin plots for both the total available CASP15 targets as well as the publicly available ones.
@@ 5.2 Molecule pretraining implies multi-ligand docking performance
Figure 6: DiffDock-L and NeuralPLexer multi-ligand predictions for CASP15 target T1187.
@@ Conclusions
In this work, we introduced PoseBench, the first deep learning benchmark for practical protein-ligand docking. Experimental results with PoseBench suggest the importance of developing new multi-ligand structure generation methods for enhanced generalization in future work. Moreover, based on these benchmark results, we posit that advances in protein-ligand docking will also likely be driven by advances in modeling macromolecular structures (Abramson et al., 2024) (e.g., by training models on full protein-nucleic acid-ligand complexes), in contrast to current methods that are trained primarily on one type of biomolecular complex (e.g., protein-ligand complexes). Key limitations of this study include its reliance on the accuracy of its predicted protein structures, its limited number of multi-ligand prediction targets available for benchmarking, and its inclusion of only a subset of all available protein-ligand docking baselines to focus on the most recent deep learning algorithms designed specifically for docking. In future work, we aim to expand not only the number of baseline methods but also the number of available multi-ligand targets while maintaining a diverse composition of heterogeneous (ionic) complexes. As a publicly available resource, PoseBench is flexible to accommodate new datasets and methods for protein-ligand structure generation.
@@
Availability. The PoseBench codebase, documentation, and tutorial notebooks are available on GitHub under a permissive MIT license, with further licensing discussed in Appendix A.
Acknowledgments and Disclosure of Funding
@@
This work was supported by a U.S. NSF grant (DBI2308699) and two U.S. NIH grants (R01GM093123 and R01GM146340). Additionally, this work was performed using high performance computing infrastructure provided by Research Support Services at the University of Missouri-Columbia (DOI: 10.32469/10355/97710). We would like to thank Martin Buttenschoen and Andriy Kryshtafovych for graciously allowing us to use in Figure 1 the PoseBusters and CASP logos, respectively. We would also like to specifically thank Martin Buttenschoen for their assistance in running the PoseBusters software suite for various protein-ligand complexes. Lastly, we would like to thank Arian Jamasb for providing insightful feedback on an early version of this manuscript.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] In this work, we have introduced PoseBench for comprehensive benchmarking of protein-ligand structure generation methods. The PoseBench benchmark results we present in this manuscript subsequently outline important research directions for future work on deep learning for protein-ligand docking including the importance of pretraining methods on large molecule corpora and evaluating them directly on multi-ligand docking targets.
    2. (b)
      Did you describe the limitations of your work? [Yes] We discuss the limitations of this study in Section Conclusions.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] In Appendix B, we discuss the broader impacts of this work as they relate to the benefits and potential risks of accurate protein-ligand structure generation methods.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] Our newly proposed protein-ligand docking benchmark adheres to the ethics review guidelines to the best of our knowledge.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] For full reproducibility and extensibility, all associated source code, data, tutorials, documentation, and benchmark method predictions are made freely available at https://github.com/BioinfoMachineLearning/PoseBench.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] In Sections 3 and 4, we carefully list the dataset preparation and method inference steps we took to produce the results we report in Section 5. Additionally, the source code and data we have made freely available will enhance readers’ ability to understand our results.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] For all generative baseline methods included in our experiments, we report the mean and standard deviation (across three independent runs) of each of the method’s corresponding performance metrics. We plot these statistical terms for each method using Matplotlib’s bar chart function [Hunter, 2007], specifically its height and yerr arguments, respectively. For completeness, like many previous works on benchmarking biomolecular methods in machine learning [Jamasb et al., 2024], we assume normally distributed errors when constructing our result plots.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] In Appendix C, we provide comprehensive details regarding the compute resources we utilized to run the experiments presented in this work.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] Please see Appendix A.
    2. (b)
      Did you mention the license of the assets? [Yes] Please see Appendix A.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] As mentioned in Appendix A, we have made freely available on Zenodo [Morehead et al., 2024] curated, deep learning-friendly versions of the Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 protein-ligand interaction datasets for apo-to-holo protein-(multi-)ligand structure generation.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] In Appendix A, we discuss the licensing and usage of each dataset that our benchmark curates in this work. In addition, we have received permission from the creators of the PoseBusters software suite and the CASP15 organizers to reference their work in our study.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] The data curated in this study is of macromolecular nature, without any personally identifiable information associated.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
[…]
[appendix]
@@ Appendix A Availability
The PoseBench codebase and tutorial notebooks are available under an MIT license at https://github.com/BioinfoMachineLearning/PoseBench. Preprocessed datasets and benchmark method predictions are available on Zenodo [Morehead et al., 2024] under a CC-BY 4.0 license, of which the Astex Diverse and PoseBusters Benchmark datasets [Buttenschoen et al., 2024] are associated with a CC-BY 4.0 license; of which the DockGen dataset [Corso et al., 2024a] is available under an MIT license; and of which the CASP15 dataset [Robin et al., 2023], as a mixture of publicly and privately available resources, is partially licensed. In particular, 15 (4 single-ligand and 11 multi-ligand targets) of the 19 CASP15 protein-ligand complexes evaluated with PoseBench are publicly available, whereas the remaining 4 (2 single-ligand and 2 multi-ligand targets) are confidential and, for the purposes of future benchmarking and reproducibility, must be requested directly from the CASP organizers. Lastly, our use of the PoseBusters software suite for molecule validity checking is permitted under a BSD-3-Clause license.
@@ Appendix B Broader impacts
Our benchmark unifies protein-ligand structure generation datasets, methods, and tasks to enable enhanced insights into the real-world utility of such methods for accelerated drug discovery and energy research. We acknowledge the risk that, in the hands of "bad actors", such technologies may be used with harmful ends in mind. However, it is our hope that efforts in elucidating the performance of recent protein-ligand structure generation methods in various macromolecular contexts will disproportionately influence the positive societal outcomes of such research such as improved medicines and subsequent clinical outcomes as opposed to possible negative consequences such as the development of new bioweapons.
@@ Appendix C Compute resources
To produce the results presented in this work, we concurrently utilized 4 80GB NVIDIA A100 GPUs for 4 weeks in total to run inference with each baseline method three times (where applicable), where each baseline deep learning method required approximately 3-7 days of GPU compute to complete its inference runs (except for FABind which completed its inference runs in the span of several hours). Notably, due to RoseTTAFold-All-Atom’s significant storage requirements for running inference with its multiple sequence alignment databases, we utilized approximately 3 TB of solid-state storage space in total to benchmark all baseline methods. Lastly, in terms of CPU requirements, our experiments utilized approximately 192 concurrent CPU processes for AutoDock Vina inference (as an upper bound) and 128 GB of CPU RAM. Note that an additional 1-2 weeks of compute were spent performing initial (failed) versions of each experiment during PoseBench’s initial phase of development.
Appendix D Metrics
@@
In this work, we reference two key metrics in the field of structural bioinformatics: RMSD and lDDT. The RMSD between a predicted ligand 3D conformation (with atomic positions $\hat{x}_{i}$ for each of the ligand’s $n$ heavy atoms) and the ground-truth conformation ($x_{i}$) is defined as:

### v4

@@ 5 Results and discussions
In this section, we present PoseBench’s results for single and multi-ligand protein-ligand structure generation and discuss their implications for future work. Note that across all the experiments, for generative methods (or methods that use generative inputs to make their predictions), we report their performance metrics in terms of the mean and standard deviation across three independent runs of the method to gain insights into its inter-run stability and consistency. For interested readers, in Appendix C, we report the average runtime and memory usage of each baseline method to determine which methods are the most efficient for real-world docking applications.
@@
Figure 2: Astex & PoseBusters dataset results for successful single-ligand docking. RMSD $\leq$ 2 Å & PB-Valid denotes a method’s percentage of ligand structures within 2 Å of the ground-truth ligand that also pass all PoseBusters filtering.
@@
Figure 3: PoseBusters dataset results for successful single-ligand docking with relaxation. RMSD $\leq$ 2 Å & PB-Valid denotes a method’s percentage of ligand structures within 2 Å of the ground-truth ligand that also pass all PoseBusters filtering.
5.1 Training on diverse clusters supports single-ligand docking performance
@@
We begin our investigations by evaluating the performance of each baseline method for single-ligand docking using the Astex Diverse and PoseBusters Benchmark datasets. Notably, for results on the PoseBusters Benchmark dataset (and subsequent datasets), we perform an additional analysis where we apply post-prediction (fixed-protein) relaxation to each method’s generated ligand conformations using molecular dynamics simulations (Eastman & Pande, 2010), as originally proposed by Buttenschoen et al. (2024). Additionally, for interested readers, in Appendix H.1 we include DockGen benchmark results for flexible-protein relaxation as implemented by Lu et al. (2024).
As shown in Figures 2 and 3, Chai-1 and DiffDock-L (in particular, the version of DiffDock employing structural cluster training (SCT)) achieve the best overall performance across both of these single-ligand datasets in terms of its percentage of correct and valid generated ligand poses (i.e., RMSD $\leq$ 2 Å & PB-Valid). To better understand this finding, in Appendix H.1, we find an even more striking instance where ablating SCT from DiffDock leads to considerably degraded docking performance for novel single-ligand protein targets. Furthermore, in the context of DockGen benchmarking, we find that Chai-1’s performance closely matches the performance of DiffDock without SCT (both notably lower than that of DiffDock-L), suggesting that training on diverse structural clusters is particularly important for docking to novel protein pockets.
@@
Figure 4: CASP15 dataset results for successful multi-ligand docking with relaxation. RMSD $\leq$ 2 Å denotes a method’s percentage of ligand structures within 2 Å of the ground-truth ligand.
@@
Figure 5: CASP15 dataset results for multi-ligand PoseBusters validity rates with relaxation. PB-Valid denotes a method’s percentage of multi-ligand structures that pass all PoseBusters filtering.
Following structural relaxation, closely behind in performance for the more challenging PoseBusters Benchmark dataset are the DL methods RoseTTAFold-AA and NeuralPLexer. Interestingly, without relaxation, AutoDock Vina combined with DiffDock-L’s predicted binding pockets achieves the third-best performance on the PoseBusters Benchmark dataset, which suggests that (1) Chai-1 and DiffDock-L are currently the only single-ligand DL methods that present a better intrinsic understanding of biomolecular physics for docking than conventional modeling tools and (2) DiffDock-L is better at locating binding pockets than standard pocket predictors such as P2Rank. Overall, these results for the Astex Diverse and PoseBusters Benchmark datasets suggest that DL methods, combined with structural relaxation, outperform conventional methods for single-ligand docking and that training future DL methods using diverse sequence (and structure)-based clusters is a promising research direction for such docking tasks. For interested readers, in Appendix H.2, we report e.g., pocket-only PoseBusters Benchmark experiments and RMSD violin plots for both the Astex Diverse and PoseBusters Benchmark datasets, which suggest that Chai-1 and DiffDock-L primarily operate in sequence and structural representation spaces, respectively.
@@
Figure 6: DiffDock-L and NeuralPLexer multi-ligand predictions for CASP15 target T1188.
5.2 Physics-informed clash penalization improves multi-ligand docking
We now turn to investigating the performance of various deep learning and conventional methods for multi-ligand docking. In contrast to the single-ligand docking results presented in Section 5.1, in Figure 4, we see a particular DL method, NeuralPLexer, stand out from all other methods in terms its multi-ligand docking performance. To better understand the factors contributing to its success, we also report results with a version of NeuralPLexer fine-tuned without its (original) van der Waals-based inter-ligand clash loss (ILCL) function (i.e., NeuralPLexer w/o ILCL), where these (ablation) results suggest that training NeuralPLexer with physics-based clash penalties has provided it with useful knowledge for successful multi-ligand docking. In contrast, all other baseline methods appear to produce only a handful of correctly docked multi-ligand poses. To more concretely understand why, in Appendix F.2, we plot the distribution of protein-ligand interactions produced by each baseline method for the CASP15 dataset, and we find that most methods struggle to correctly capture e.g., the distribution of hydrophobic interactions or Van der Waals contacts this dataset presents. Using CASP15 target T1188 as a case study, in Figure 6, we illustrate how this distributional mismatch often leads to methods such as DiffDock-L producing top-ranked predictions with multi-ligand steric clashes that must be (unoptimally) resolved using structural relaxation. To summarize, we find that these interaction-level distribution mismatches translate to poor multi-ligand docking performance for most baseline methods and that NeuralPLexer’s inter-ligand clash loss has improved its ability to match the ground-truth distribution of CASP15 protein-ligand interactions for multi-ligand docking.
To further inspect each method’s understanding of biomolecular physics for multi-ligand docking, in Figure 5 we report each method’s percentage of predicted protein-ligand complexes (whether correct or not) for which all ligand conformations in the complex are jointly considered valid according to the PoseBusters software suite (i.e., PB-Valid). In short, in the context of multi-ligands, we find that NeuralPLexer and AutoDock Vina are nearly tied in terms of their PoseBusters validity rates following structural relaxation and that Ensemble (Con) provides the best validity rates overall. To better understand this latter result, we note that NeuralPLexer’s predictions seem to be among the most frequently selected by Ensemble (Con) for multi-ligand prediction targets (n.b., and conversely DiffDock-L for single-ligand targets), which suggests that NeuralPLexer consistently produces the highest percentage of valid ligand poses for a given multi-ligand complex, further supporting the notion that NeuralPLexer’s multi-ligand training protocol has improved its understanding of protein-ligand binding patterns crucial for multi-ligand docking. For interested readers, in Appendix H.3, we report additional results e.g., in terms of lDDT-PLI and RMSD violin plots for both the total available CASP15 targets as well as those publicly available.
@@ Conclusions
In this work, we introduced PoseBench, the first deep learning (DL) benchmark for broadly applicable protein-ligand docking. Benchmark results with PoseBench currently suggest a negative answer to the question ”Are we there yet (for structural drug discovery) with DL-based protein-ligand docking?”. In this work, we have observed that while DL methods such as Chai-1 and DiffDock-L can identify the correct binding pockets in many single-ligand protein targets, most DL methods struggle to generalize to multi-ligand docking targets. Based on these results, for the development of future DL docking methods, we recommend researchers train new docking methods directly (1) on structurally clustered multi-ligand protein complexes available in new DL-ready biomolecular datasets (Abramson et al., 2024; Wang & Morehead, 2024) (2) using physics-informed inter-ligand steric clash penalties (Qiao et al., 2024). Key limitations of this study include its reliance on the accuracy of its predicted protein structures, its (currently) limited number of multi-ligand prediction targets available for benchmarking, and its inclusion of only a subset of all available protein-ligand docking baselines to focus on the most recent deep learning algorithms designed specifically for docking and structure generation. In future work, we aim to expand not only the number of baseline methods but also the number of available (CASP) multi-ligand targets while maintaining a diverse composition of heterogeneous (ionic) complexes. As a publicly available resource, PoseBench is flexible to accommodate new datasets and methods for protein-ligand structure generation.
@@
Availability. The PoseBench codebase, documentation, and tutorial notebooks are available at https://github.com/BioinfoMachineLearning/PoseBench under a permissive MIT license, with further licensing discussed in Appendix A.
Author Contributions
AM designed the experiments in this work, wrote their supporting code, analyzed their results, and described them in the current manuscript. NG and JL helped run the conventional docking baseline methods for the experiments in this work and helped revise the current manuscript. JC provided funding, support, important feedback, and manuscript revisions for this work.
Acknowledgments
@@
This work was supported by a U.S. NSF grant (DBI2308699) and two U.S. NIH grants (R01GM093123 and R01GM146340). Additionally, this work was performed using high performance computing infrastructure provided by Research Support Services at the University of Missouri-Columbia (DOI: 10.32469/10355/97710). We would like to thank Martin Buttenschoen and Andriy Kryshtafovych for graciously allowing us to use in Figure 1 the PoseBusters and CASP logos, respectively. We would also like to specifically thank Martin Buttenschoen for their assistance in running the PoseBusters software suite for various protein-ligand complexes and Pawan Neupane for their help in running different structure prediction models. Lastly, we would like to thank Arian Jamasb for providing insightful feedback on an early version of this manuscript.
@@ Appendix A Availability
The PoseBench codebase and tutorial notebooks are available under an MIT license at https://github.com/BioinfoMachineLearning/PoseBench. Preprocessed datasets and benchmark method predictions are available on Zenodo (Morehead et al., 2024) under a CC-BY 4.0 license, of which the Astex Diverse and PoseBusters Benchmark datasets (Buttenschoen et al., 2024) are associated with a CC-BY 4.0 license; of which the DockGen dataset (Corso et al., 2024a) is available under an MIT license; and of which the CASP15 dataset (Robin et al., 2023), as a mixture of publicly and privately available resources, is partially licensed. In particular, 15 (4 single-ligand and 11 multi-ligand targets) of the 19 CASP15 protein-ligand complexes evaluated with PoseBench are publicly available, whereas the remaining 4 (2 single-ligand and 2 multi-ligand targets) are confidential and, for the purposes of future benchmarking and reproducibility, must be requested directly from the CASP organizers. Notably, the pre-holo-aligned protein structures predicted by AlphaFold 3 for these four benchmark datasets (available on Zenodo (Morehead et al., 2024)) must only be used in accordance with the Terms of Service provided by the AlphaFold Server. Lastly, our use of the PoseBusters software suite for molecule validity checking is permitted under a BSD-3-Clause license.
@@ Appendix B Broader impacts
Our benchmark unifies protein-ligand structure generation datasets, methods, and tasks to enable enhanced insights into the real-world utility of such methods for accelerated drug discovery and energy research. We acknowledge the risk that, in the hands of ”bad actors”, such technologies may be used with harmful ends in mind. However, it is our hope that efforts in elucidating the performance of recent protein-ligand structure generation methods in various macromolecular contexts will disproportionately influence the positive societal outcomes of such research such as improved medicines and subsequent clinical outcomes as opposed to possible negative consequences such as the development of new bioweapons.
@@ Appendix C Compute resources
To produce the results presented in this work, we ran a high performance computing sweep that concurrently utilized 24 80GB NVIDIA A100 GPU nodes for 3 days in total to run inference with each baseline method three times (where applicable), where each baseline deep learning (DL) method required approximately 8 hours of GPU compute to complete its inference runs (except for FABind which completed its inference runs in the span of a couple hours). Notably, due to RoseTTAFold-All-Atom’s significant storage requirements for running inference with its multiple sequence alignment databases, we utilized approximately 3 TB of solid-state storage space in total to benchmark all baseline methods. Lastly, in terms of CPU requirements, our experiments utilized approximately 64 concurrent CPU threads for AutoDock Vina inference (as an upper bound) and 60 GB of CPU RAM. Note that an additional 4-5 weeks of compute were spent performing initial (non-sweep) versions of each experiment during PoseBench’s initial phase of development.
@@
Appendix D Metrics
In this work, we reference two key metrics in the field of structural bioinformatics: RMSD and lDDT. The RMSD between a predicted 3D conformation (with atomic positions $\hat{x}_{i}$ for each of the molecule’s $n$ heavy atoms) and the ground-truth conformation ($x_{i}$) is defined as: